Fred purchased 100 shares of Ekto Corporation common stock for $30.3 per share on June 1,...

  1. Fred purchased 100 shares of Ekto Corporation common stock for $30.3 per share on June 1, 2001 and then sold it exactly 11 years later for $43.91 per share. What was Fred's geometric mean average annual return on this investment?

    4.08%

    2.82%

    3.43%

    2.48%

1 points   

QUESTION 6

  1. Fred purchased 100 shares of Ekto Corporation common stock for $31.99 per share on June 1, 2001 and then sold it exactly 3 years later for $43.53 per share. What was Fred's geometric mean average annual return on this investment?

    12.02%

    8.84%

    8.15%

    10.81%
Answer #1

Geometric Mean Average Annual Return = Compounded Annual Growth Rate (CAGR)

CAGR = (Future Value/Initial Investment)1/number of periods -1

5)

Geometric Mean = (43.91/30.3)1/11 -1 = 1.44917490.090909 -1 = 1.0343 - 1 = 0.0343 = 3.43%

6)

Geometric Mean = (43.53/31.99)1/3 - 1 = 1.36073770.333333 -1 = 1.10813 -1 = 0.10813 = 10.813%

(If this was helpful then please rate positively. Thank You:)